OpenStudy (anonymous):

Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (mimi_x3):

Elimination method.

OpenStudy (cathyangs):

I would use elimination, because all the coefficients have no common factor bigger than 1, and the x coefficients are additive inverses. This will make it easier to use elimination. 2x+ 5y=-3 +-15x-5y=-10 -------------- -13x =-13 using algebra: x=1. 2(1)+5y=-3 2+5y=-3 5y=-5 y=-1 (x,y) = (1,-1)